Ms. Thompson has written a log for her store every 3 months for the last 6 years. If 2/3 of the log shows the store earns more money than it spends.

1. How many logs show her store spending money than it earns?

Found 2 solutions by ikleyn, josmiceli:
Answer by ikleyn(53763)

.
Ms. Thompson has written a log for her store every 3 months for the last 6 years. If 2/3 of the log shows the store earns more money than it spends.
1. How many logs show her store spending money than it earns?
~~~~~~~~~~~~~~


Read your question, written under N1, and try to understand what is written there.


When you post  GIBBERISH, you  steal  my time  (!)



Answer by josmiceli(19441)

( 12 months/year ) x ( 6 years ) = ( 72 months )
and
( 72 months ) / ( 3 months/ log ) = ( 24 logs )
and
( 24 logs ) x ( 2/3 ) = ( 16 logs ) earns more than spends
and
( 24 logs ) x ( 1/3 ) = ( 8 logs ) spends more than earns